# Towards Graph Foundation Models: A Study on the Generalization of Positional and Structural Encodings

### v1

@@ 5 Experimental Results
Overall, our experiments include three dataset types: (i) synthetic datasets EXP, CEXP, CSL, TRI, and TRIX [11, 49, 12], designated to study the expressivity of GPSE, as reported in Table 1. (ii) ZINC-12k, where we find GPSE to perform well across all considered aspects, and (iii) MolNet, where we find that GPSE does not perform better than baseline methods in general. The evaluations on these diverse datasets allow us to provide well-grounded insights regarding the utilization of GPSE as a component in future GFMs. We provide details on the learning and evaluation procedures in Appendix F. We elaborate on the considered PSEs in Appendix E. We elaborate on the datasets and the motivation for using them in Appendix G.
@@ 5.1 GPSEs Convergence to PSEs Performance
To evaluate the speedup GPSE enables, we compare how many epochs a GPSE-augmented GNN takes to reach the same performance as other PSE-augmented GNNs, as shown in Table 2. Our results show that using GPSE yields good performance significantly faster compared with all other PSEs. The smallest speedup is a factor of 2, and in almost all cases, GPSE obtains a speedup factor of 10 or more. Reaching good performance faster not only saves training time during parameter tuning but also reduces the environmental impact by lowering CO2 emissions associated with prolonged computational efforts. We note here, that Table 2 does not imply that GPSE converges to its optimal performance faster, it only implies that GPSE attains good performances faster than other PSEs. In fact, according to our evaluations, GPSE converges at the same speed as other PSEs (see Appendix H).
@@ 5.2 Influence of Sample Size
A common challenge in machine learning is the scarcity of available training data, which can significantly hinder model performance. It is common in areas like NLP and CV to finetune pre-trained models in data-scarce applications. Some prominent examples are LLMs like GPT-3, which can be fine-tuned or used in zero-shot settings for various NLP tasks [23]. In CV, models like CLIP, which learn visual concepts from natural language supervision, have been effectively adapted to new tasks with minimal data [55]. These approaches leverage powerful representations learned from large-scale data, making them highly effective in transfer learning scenarios.
To investigate how pre-trained GPSE models can mitigate the impact of limited data, we conducted a few-shot learning-inspired experiment. In this experiment, we progressively reduce the amount of training data for the downstream task, and evaluate the resulting performance of a downstream GNN when combined with different PSEs against GPSE and its variants. As proposed in an ablation study in Cantürk et al. [21], we start with all training data ( training, validation, testing) and iteratively halve the training data to just , while leaving the validation and test data unchanged.
Fig. 2 shows the results of the experiment on ZINC-12k and MolNet ToxCast datasets. GPSE’s performance is used as a baseline (black dotted line) and compared with other PSEs, including LapPE, RWSE, and the GPSE-and GPSE+variants. Both GPSE-and GPSE+perform similarly to GPSE. On ZINC-12k, GPSE-like models outperform all other PSEs, with the gap widening as data decreases logarithmically. However, this advantage is not consistent across all datasets. For instance, on ToxCast, GPSE-like models outperform other PSEs with full training data but are surpassed by other PSEs in data-scarce scenarios. Additional results are provided in Fig. 3 (Appendix I), highlighting the significant benefits of learned PSEs in future GFMs.
@@ 5.3 Generalization of PSEs
We now investigate why GPSE might or might not outperform other PSEs when used to augment downstream GNN models. We focus on two possible proxies for good performance: (i) better training loss and (ii) better generalization error. Notably, if one is kept constant, then the respective other will govern changes in test performance. In this set of experiments, we define the generalization error as the difference between train performance and test performance, where the performance metric depends on the dataset and task of interest. For example, in the case of regression tasks such as on ZINC-12k, a larger value—based on MAE—means better generalization, while for classification tasks, a lower value—based on AUROC–indicates better generalization. We provide the obtained generalization errors for ZINC-12k in Table 3, and the respective train (and test) mean-absolute-error (MAE) in Table 5 (and Table 16 in Appendix J). On ZINC-12k the GPSE variants attain the best test performance regardless of the backbone (Table 16 in Appendix). Table 3 shows the GPSE variants attain the best generalization errors compared with other PSEs. In contrast, Table 5 shows that while GPSE attains fairly low training MAE, other PSEs “outperform” the GPSE variants in terms of training performance. These observations imply that the performance gains of GPSE-augmented models are primarily due to better generalization, and not, for instance, due to a model that better fits the training set. However, this better generalization does not always hold. In contrast to previous observations, Table 4 showing the generalization error for the MolNet datasets, demonstrates that GPSE and its variants can be outperformed considering generalization on some downstream tasks (MUV and BACE are notable examples). Better generalization and generally better performance are desired properties of a foundation model. These results indicate that while GPSE shows some superior generalization capabilities, they need to be examined per case. While this work studied expressivity in Section 4 and generalization here, we are not attempting to conflate the two herein. The observations on expressivity and generalization are entirely separate. Notably, even though GPSE-, GPSE, and GPSE+have different practical expressivity, their generalization seems comparable which agrees with previous work on the complexity of expressivity and generalization [56].
@@ 6 Discussion and Conclusion
Future research should enhance generalization in GNNs, especially for graph SEs and PEs. Initial studies on the generalization of PSEs and embeddings for substructures, such as [56], suggest that subgraph-enhanced WL kernels’ generalization depends on the counted substructures. This may apply to MPNNs as well; for instance, PSEs used in GPSE pre-training generalize well on ZINC but poorly on some MolNet datasets. The broader impact of graph structure on generalization has been identified as a key challenge in ML on graphs (Challenge 3.2 in Morris et al. [31]), and addressing these challenges will be crucial for advancing GFMs.
@@ Acknowledgements
Part of this work was conducted within the DFG Research Unit FOR 5359 on Deep Learning on Sparse Chemical Process Data (KL 2698/6-1 and KL 2698/7-1). The authors acknowledge support by the Carl-Zeiss Foundation (AI-Care, Process Engineering 4.0), the DFG awards INST 248/377 (DFG TRR 375 ), KL 2698/5-2, FE 2282/6-1, and FE 2282/1-2, and the BMBF award 01—S2407A.
@@ Appendix A Expressivity
For the most part, in this work, we only consider expressivity attained from RNF because it was proposed as the method of attaining expressivity in Cantürk et al. [21]. Here we now discuss the reason for this choice as well as some related work on the topic of expressivity in machine learning on graphs. Related work on expressivity has focussed on two main directions for attaining expressivity for MPNNs.
@@
The first direction, which we discussed in detail in this paper, is essentially the use of PSEs. Notably PSEs like LapPE [16, 19] and RWSE [19, 15, 38] increase expressivity when used to augment downstream MPNNs but have not been shown to be universal and in fact we show in this work that they cannot be universal. Further PSEs that are well-known to be universal are based on subgraph counting [37, 41, 40, 42] or homomorphism counting [57, 41, 47]. However, related work usually considers the setting of uncolored graphs, in which case these PSE are indeed universal if sufficiently many subgraphs/homomorphisms are being counted. Once more, according to Theorem 2 subgraph and homomorphism counting that disregards node colors will not grant universality to augmented downstream MPNNs considering colored graphs. The emphasis here is on PSEs that disregard node colors, one could propose subgraph and homomorphism counting of colored subgraphs which would be fully universal. Lastly, RNF which can be considered a PSE has been shown to grant universality when used to augment downstream MPNNs [11, 12, 44]. As RNF relies on randomness to differentiate nodes during computation, RNF is universal regardless of node colors. The universality or expressivity of PSEs is inherently linked to the individualization refinement algorithm for which PSEs act as individualization functions as discussed in [44] which is inherently efficient assuming that any precomputations of the PSEs can be done efficiently.
@@
The second direction modifies an MPNN’s computation to attain more expressivity. Common examples include the k-GNN [3, 58, 59] and PPGN [45, 60]. These higher-order methods are essentially related to the -dimensional Weisfeiler-Leman algorithm and, as a result, suffer from high computational and memory costs. However, they notably show advantages in performance in some cases.
@@
GPSE [21] used RNF instead of other methods of expressivity. Notably, of the methods mentioned, only RNF, subgraph, and homomorphism counting, and higher-order methods are truly universal. Subgraph and homomorphism counting, however, require a potentially very large set of graphs to be counted (which is even bigger if not infinitely large for colored graphs), and higher-order methods might require a very large order (in the case of k-GNNs, a very large ) for which the computational and memory cost can be prohibitively large. In this sense, then, RNF is the only simple, effective, and scalable method of granting GPSE universality, which makes it the best choice for GPSE.
Appendix B Further Details on Graph Positional and Structural Encoder
GPSE [21] is a recently introduced GNN model that aims to learn and predict robust and transferable representations for positional and structural encodings using RNFs as inputs, and constitutes the main architecture studied in our paper. In this paper, we choose to work with GPSE because it generalizes other PSEs, such as the graph Laplacian eigenvectors or random-walk structural encodings, in the sense that it learns these PSEs during its training phase. In what follows, we provide a brief standalone introduction to the model; further information is available via the original paper.
GPSE model GPSE employs an encoder-decoder architecture in which the encoder and decoder is trained jointly, whereas in inference only the encoder is used. The encoder is a deep MPNN consisting of stacked GatedGCN [53] layers, with residual gating and skip-connections in-between layers. The decoder consists of multiple node-level MLP heads, each corresponding to a different PSE component.
with and are learnable parameters.
GPSE training and usage Cantürk et al. [21] train GPSE by minimizing the sum of the reconstruction losses for each target PSE (details on PSEs in Appendix E), where the reconstruction loss for a given PSE is defined as the sum of loss and cosine similarity losses between the predicted and target values, summed across all nodes. Thus for target PSEs, the loss for a given graph becomes:
where and denote the true and predicted values for node for the th PSE, while and denote their respective normalized versions.
GPSE is designed to be used primarily as a pre-trained model, such that after pre-training on a large and diverse set of graphs, it can reliably and efficiently extract PSE representations for unseen graphs by using the encoder in inference mode. The authors pre-train GPSE on MolPCBA [61] due to the higher number of unique graph structures compared to larger datasets like PCQM4Mv2 [62] as original node features are not considered in GPSE training. Using this MolPCBA-pre-trained GPSE model, the authors predict PSE representations for a large variety of datasets drawn from a variety of domains, and demonstrate that these GPSE encodings are either significantly more effective than, or just as effective as conventional PSEs in the worst case even for significantly OOD graphs.
Appendix C Proof of Theorem 1
@@ A GIN layer is defined as follows:
Assuming that is bounded, this quantity can be arbitrarily small by choosing small enough. To be clear, the assumption on the input we require, is that is bounded and that each node caries a feature that is constant over all nodes, in this proof we assumed this constant to be 1, however, this is true for any such non-zero constant. Notably, we can also make weaker assumptions, by assuming that there is a linear combination of features that is constant over all nodes. However, in practice, we can simply add a constant 1 as a feature to each node, which is a common augmentation to ensure MPNNs can count the size of neighborhoods. ∎
Appendix D Additional expressivity experiment
@@
Here we include an additional expressivity experiment based on Huang et al. [63]. The task is to count the inclusion of nodes in certain substructures on random Erdős-Rényi graphs. As in the original work, we train 5-layer GatedGNNs [64]. We run the experiment with 10 different random seeds and report the test MAE divided by the label standard deviation referred to as normalized MAE in Table 6 and Table 7. Since GPSE is trained to predict cycle counts we would expect GPSE to perform well in Table 6. However, as demonstrated previously, GPSE only learns to count cycles for molecule-like graphs, because these are the only graphs included in the pre-training set. GPSE+has additional hard graphs added to its pre-training set which appears to help its performance somewhat, although its performance is still a far cry from other methods reported in Huang et al. [63]. Perhaps adding Erdős-Rényi graphs instead of random regular graphs would help with its performance. In some instances GPSE-performs best, which might be due to GPSE-overfitting less to the use of RNI.
Appendix E PSEs
@@
The absolute value operation is needed to counter the sign ambiguity of the graph Laplacian eigenvectors, a known issue to many previous works that use the Laplacian eigenvectors to augment the models [65, 18]. We maintain learning this absolute value function of the eigenvectors as per Cantürk et al. [21], and similarly learn the eigenvalues as graph-level attributes.
@@
The RWSE [38] is defined as the probability of returning back to the starting state of a random walk after steps of a random walk:
@@
ElstaticPE for any node is then a collection of statistics that summarizes the electrostatic interaction of with other nodes, such as the minimum, mean and standard deviations of potentials from to all other nodes or only its neighbors [16]. We defer to Cantürk et al. [21] for more detailed definitions of its individual components.
@@
Appendix F Training and Evaluation Details
@@
MoleculeNet small benchmarks settings We used the default GINE architecture following previous studies [66], which has five hidden layers and 300 hidden dimensions. For all five benchmarks, we use the same GPSE processing encoder settings as shown in Table 9(a).
CSL, EXP, and TRI synthetic graph benchmarks settings We follow He et al. [67] and use GIN [8] as the underlying MPNN model, with five hidden layers and 128 dimensions. We use the same GPSE processing encoder settings for CSL, EXP, CEXP, TRI, and TRIX as shown in Table 9(b).
@@
F.1 Convergence to PSEs Performance

### v2

@@ 5 Experimental Results
Overall, our experiments include three dataset types: (i) synthetic datasets EXP, CEXP, CSL, TRI, and TRIX (Abboud et al., 2021; Murphy et al., 2019; Sato et al., 2021), designated to study the expressivity of GPSE, as reported in Table 1. (ii) ZINC-12k, where we find GPSE to perform well across all considered aspects, and (iii) MolNet, where we find that GPSE does not perform better than baseline methods in general. The evaluations on these diverse datasets allow us to provide well-grounded insights regarding the utilization of GPSE as a component in future GFMs. We provide details on the learning and evaluation procedures in Appendix E. We elaborate on the considered PSEs in Appendix D. We elaborate on the datasets and the motivation for using them in Appendix F.
@@ 5.1 GPSEs Convergence to PSEs Performance
To evaluate the speedup GPSE enables, we compare how many epochs a GPSE-augmented GNN takes to reach the same performance as other PSE-augmented GNNs, as shown in Table 2. Our results show that using GPSE yields good performance significantly faster compared with all other PSEs. The smallest speedup is a factor of 2, and in almost all cases, GPSE obtains a speedup factor of 10 or more. Reaching good performance faster not only saves training time during parameter tuning but also reduces the environmental impact by lowering CO2 emissions associated with prolonged computational efforts. We note here, that Table 2 does not imply that GPSE converges to its optimal performance faster, it only implies that GPSE attains good performances faster than other PSEs. In fact, according to our evaluations, GPSE converges at the same speed as other PSEs (see Appendix G).
@@ 5.2 Influence of Sample Size
A common challenge in machine learning is the scarcity of available training data, which can significantly hinder model performance. It is common in areas like NLP and CV to finetune pre-trained models in data-scarce applications. Some prominent examples are LLMs like GPT-3, which can be fine-tuned or used in zero-shot settings for various NLP tasks (Brown et al., 2020). In CV, models like CLIP, which learn visual concepts from natural language supervision, have been effectively adapted to new tasks with minimal data (Radford et al., 2021). These approaches leverage powerful representations learned from large-scale data, making them highly effective in transfer learning scenarios.
To investigate how pre-trained GPSE models can mitigate the impact of limited data, we conducted a few-shot learning-inspired experiment. In this experiment, we progressively reduce the amount of training data for the downstream task, and evaluate the resulting performance of a downstream GNN when combined with different PSEs against GPSE and its variants. As proposed in an ablation study in Cantürk et al. (2024), we start with all training data ( training, validation, testing) and iteratively halve the training data to just , while leaving the validation and test data unchanged.
Fig. 2 shows the results of the experiment on ZINC-12k and MolNet ToxCast datasets. GPSE’s performance is used as a baseline (black dotted line) and compared with other PSEs, including LapPE, RWSE, and the GPSE-and GPSE+variants. Both GPSE-and GPSE+perform similarly to GPSE. On ZINC-12k, GPSE-like models outperform all other PSEs, with the gap widening as data decreases logarithmically. However, this advantage is not consistent across all datasets. For instance, on ToxCast, GPSE-like models outperform other PSEs with full training data but are surpassed by other PSEs in data-scarce scenarios. Additional results are provided in Fig. 3 (Appendix H), highlighting the significant benefits of learned PSEs in future GFMs.
@@ 5.3 Generalization of PSEs
We now investigate why GPSE might or might not outperform other PSEs when used to augment downstream GNN models. We focus on two possible proxies for good performance: (i) better training loss and (ii) better generalization error. Notably, if one is kept constant, then the respective other will govern changes in test performance. In this set of experiments, we define the generalization error as the difference between train performance and test performance, where the performance metric depends on the dataset and task of interest. For example, in the case of regression tasks such as on ZINC-12k, a larger value—based on MAE—means better generalization, while for classification tasks, a lower value—based on AUROC–indicates better generalization. We provide the obtained generalization errors for ZINC-12k in Table 3, and the respective train (and test) mean-absolute-error (MAE) in Table 5 (and Table 16 in Appendix I). On ZINC-12k the GPSE variants attain the best test performance regardless of the backbone (Table 16 in Appendix). Table 3 shows the GPSE variants attain the best generalization errors compared with other PSEs. In contrast, Table 5 shows that while GPSE attains fairly low training MAE, other PSEs “outperform” the GPSE variants in terms of training performance. These observations imply that the performance gains of GPSE-augmented models are primarily due to better generalization, and not, for instance, due to a model that better fits the training set. However, this better generalization does not always hold. In contrast to previous observations, Table 4 showing the generalization error for the MolNet datasets, demonstrates that GPSE and its variants can be outperformed considering generalization on some downstream tasks (MUV and BACE are notable examples). Better generalization and generally better performance are desired properties of a foundation model. These results indicate that while GPSE shows some superior generalization capabilities, they need to be examined per case. While this work studied expressivity in Section 4 and generalization here, we are not attempting to conflate the two herein. The observations on expressivity and generalization are entirely separate. Notably, even though GPSE-, GPSE, and GPSE+have different practical expressivity, their generalization seems comparable, which agrees with previous work on the complexity of expressivity and generalization (Franks et al., 2024).
@@ 6 Discussion and Conclusion
Future research should enhance generalization in GNNs, especially for graph SEs and PEs. Initial studies on the generalization of PSEs and embeddings for substructures, such as in Franks et al. (2024), suggest that subgraph-enhanced WL kernels’ generalization depends on the counted substructures. This may apply to MPNNs as well; for instance, PSEs used in GPSE pre-training generalize well on ZINC but poorly on some MolNet datasets. The broader impact of graph structure on generalization has been identified as a key challenge in ML on graphs (Challenge 3.2 in Morris et al. (2024)), and addressing these challenges will be crucial for advancing GFMs.
@@ Acknowledgements
Part of this work was conducted within the DFG Research Unit FOR 5359 on Deep Learning on Sparse Chemical Process Data (BU 4042/2-1, KL 2698/6-1, and KL 2698/7-1) and the DFG TRR 375 (no. 511263698). SF and MK acknowledge support by the Carl-Zeiss Foundation (AI-Care, Process Engineering 4.0), the BMBF award 01|S2407A, and the DFG awards FE 2282/6-1, FE 2282/1-2, KL 2698/5-2, and KL 2698/11-1. ME is funded by the Blavatnik-Cambridge fellowship, the Cambridge Accelerate Programme for Scientific Discovery and the Maths4DL EPSRC Programme. Additional support is by Bourse en intelligence artificielle des Études supérieures et postdoctorales (ESP) 2023-2024 [SC]; Canada CIFAR AI Chair, IVADO (Institut de valorisation des données) grant PRF-2019-3583139727, FRQNT (Fonds de recherche du Québec - Nature et technologies) grant 299376, NSERC Discovery grant 03267 and NSF DMS grant 2327211 [GW].
@@ Appendix A Expressivity
For the most part, in this work, we only consider expressivity attained from RNF because it was proposed as the method of attaining expressivity in Cantürk et al. [2024]. Here we now discuss the reason for this choice as well as some related work on the topic of expressivity in machine learning on graphs. Related work on expressivity has focused on two main directions for attaining expressivity for MPNNs.
@@
The first direction, which we discussed in detail in this paper, is essentially the use of PSEs. Notably PSEs like LapPE [Kreuzer et al., 2021, Rampášek et al., 2022] and RWSE [Rampášek et al., 2022, Dwivedi and Bresson, 2021, Dwivedi et al., 2022b] increase expressivity when used to augment downstream MPNNs but have not been shown to be universal and in fact we show in this work that they cannot be universal. Further PSEs that are well-known to be universal are based on subgraph counting [Shiv and Quirk, 2019, Bouritsas et al., 2022, Zhao et al., 2022, Chen et al., 2022] or homomorphism counting [Jin et al., 2024, Bouritsas et al., 2022, Zhang et al., 2024]. However, related work usually considers the setting of uncolored graphs, in which case these PSE are indeed universal if sufficiently many subgraphs/homomorphisms are being counted. Once more, according to Theorem 2 subgraph and homomorphism counting that disregards node colors will not grant universality to augmented downstream MPNNs considering colored graphs. The emphasis here is on PSEs that disregard node colors, one could propose subgraph and homomorphism counting of colored subgraphs which would be fully universal. Lastly, RNF which can be considered a PSE has been shown to grant universality when used to augment downstream MPNNs [Abboud et al., 2021, Sato et al., 2021, Franks et al., 2023]. As RNF relies on randomness to differentiate nodes during computation, RNF is universal regardless of node colors. The universality or expressivity of PSEs is inherently linked to the individualization refinement algorithm for which PSEs act as individualization functions as discussed in Franks et al. [2023], which is inherently efficient assuming that any precomputations of the PSEs can be done efficiently.
@@
The second direction modifies an MPNN’s computation to attain more expressivity. Common examples include the k-GNN [Morris et al., 2019, 2020, Wang et al., 2022b] and PPGN [Maron et al., 2019a, b]. These higher-order methods are essentially related to the -dimensional Weisfeiler-Leman algorithm and, as a result, suffer from high computational and memory costs. However, they notably show advantages in performance in some cases.
@@
GPSE [Cantürk et al., 2024] used RNF instead of other methods of expressivity. Notably, of the methods mentioned, only RNF, subgraph, and homomorphism counting, and higher-order methods are truly universal. Subgraph and homomorphism counting, however, require a potentially very large set of graphs to be counted (which is even bigger if not infinitely large for colored graphs), and higher-order methods might require a very large order (in the case of k-GNNs, a very large ) for which the computational and memory cost can be prohibitively large. In this sense, then, RNF is the only simple, effective, and scalable method of granting GPSE universality, which makes it the best choice for GPSE.
Lastly, in this work we focus mostly on expressivity as it pertains to the 1-WL, however, some more finegrained approaches to expressivity have been recently proposed. For example, expressivity based on biconnectivity [Zhang et al., 2023] or expressivity based on homomorphism counting [Zhang et al., 2024].
Appendix B Proof of Theorem 1
@@ A GIN layer is defined as follows:
Assuming that is bounded, this quantity can be arbitrarily small by choosing small enough. To be clear, the assumption on the input we require is that is bounded and that each node carries a feature that is constant over all nodes, in this proof we assumed this constant to be 1, however, this is true for any such non-zero constant. Notably, we can also make weaker assumptions, by assuming that there is a linear combination of features that is constant over all nodes. However, in practice, we can simply add a constant 1 as a feature to each node, which is a common augmentation to ensure MPNNs can count the size of neighborhoods. ∎
Appendix C Additional expressivity experiment
@@
Here we include an additional expressivity experiment based on Huang et al. [2023b]. The task is to count the inclusion of nodes in certain substructures on random Erdős-Rényi graphs. As in the original work, we train 5-layer GatedGNNs [Li et al., 2013]. We run the experiment with 10 different random seeds and report the test MAE divided by the label standard deviation referred to as normalized MAE in Table 6 and Table 7. Since GPSE is trained to predict cycle counts, we would expect GPSE to perform well in Table 6. However, as demonstrated previously, GPSE only learns to count cycles for molecule-like graphs, because these are the only graphs included in the pre-training set. GPSE+has additional hard graphs added to its pre-training set, which appears to help its performance somewhat, although its performance is still a far cry from other methods reported in Huang et al. [2023b]. Perhaps adding Erdős-Rényi graphs instead of random regular graphs would help with its performance. In some instances GPSE-performs best, which might be due to GPSE-overfitting less to the use of RNI.
Appendix D PSEs
@@
The absolute value operation is needed to counter the sign ambiguity of the graph Laplacian eigenvectors, a known issue to many previous works that use the Laplacian eigenvectors to augment the models [Dwivedi et al., 2020, Lim et al., 2023]. We maintain learning this absolute value function of the eigenvectors as per Cantürk et al. [2024], and similarly learn the eigenvalues as graph-level attributes.
@@
The RWSE [Dwivedi et al., 2022b] is defined as the probability of returning back to the starting state of a random walk after steps of a random walk:
@@
ElstaticPE for any node is then a collection of statistics that summarizes the electrostatic interaction of with other nodes, such as the minimum, mean and standard deviations of potentials from to all other nodes or only its neighbors [Kreuzer et al., 2021]. We defer to Cantürk et al. [2024] for more detailed definitions of its individual components.
@@
Appendix E Training and Evaluation Details
@@
MoleculeNet small benchmarks settings We used the default GINE architecture following previous studies [Hu* et al., 2020], which has five hidden layers and 300 hidden dimensions. For all five benchmarks, we use the same GPSE processing encoder settings as shown in Table 9(a).
CSL, EXP, and TRI synthetic graph benchmarks settings We follow He et al. [2023] and use GIN [Xu et al., 2019] as the underlying MPNN model, with five hidden layers and 128 dimensions. We use the same GPSE processing encoder settings for CSL, EXP, CEXP, TRI, and TRIX as shown in Table 9(b).
@@
E.1 Convergence to PSEs Performance